the coconut oil demand function is a=1200 -9.5p + 16.2p_p + 0.2y , where Q is the quantity of coconut oil in cents per pound , P_p is the price of palm oil in cents per pound and y is the income of consumers. assume that p is initially 45 cents per pound, pp is 31 cents per pound and q is 1275 thousand metric tons per year. calculate the income elasticity of demand for coconut oil


1
Expert's answer
2021-03-02T07:57:05-0500

"Q=1200-9.5P+16.2Pp+0.2Y"

where P= 45c or 0.45

Pp=31c or 0.31

Q=1275

Y= "\frac{1275-(1200-9.5(0.45)+16.2(0.31)}{0.2}"

= 371.2656c ( 371.3c) or 3.71

"Ey=\frac{\delta{Q}}{\delta{Y}}(\frac{Y}{Q})"

Where "\frac{\delta{Q}}{\delta{Y}}=0.2"

Ey=0.2"(\frac{371.3}{1275})"

Ey =0.2(0.2912)

Ey=0.058